### v2

@@ 6 Conclusion
Although using an ensemble of models can provide a powerful deferral rule for cascading, the additional costs required to compute such an ensemble may not lead to savings in all inference scenarios. Despite this, our work shows that this simple approach is surprisingly effective given the large differences in model sizes that reach state-of-the-art accuracy in recent ML tasks. We demonstrate improvements via a number of real-world case studies, including a study on communication costs in edge-to-cloud inference, rental costs in cloud-based settings, and the cost of black-box API services. Overall, our results demonstrate ABC’s capacity to improve the efficiency of adaptive inference systems without the complexities associated with traditional cascade frameworks, making it a compelling option for practitioners focused on reducing inference latency. In future work, we aim to extend the approach herein to open generation tasks.
7 Acknowledgements
We thank Xiaofang Wang, Wittawat Jitkrittum, Lucio Dery, Pratiksha Thaker, and Kevin Kuo for their helpful comments. This work was supported in part by the National Science Foundation grants IIS2145670, CCF2107024, IIS1705121, IIS1838017, IIS2046613, IIS2112471, and funding from Amazon, Apple, Google, Intel, Meta, Morgan Stanley, Scribe, and the CyLab Security and Privacy Institute. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any of these funding agencies.
@@ Appendix B Estimating Agreement Threshold $\theta$
Figure 6: Estimation of agreement threshold $\theta$ stability as a function of the number of samples used, across different model accuracy levels on the ImageNet1k dataset. Each plot corresponds to a model with a specific accuracy, shown in the legend. The initial estimate is with $100$ samples, with subsequent estimates with larger and larger number of samples. The initial estimate is reasonably close within subsequent estimates with larger number of samples.
@@ Appendix C Existence of Safe Deferral Rules and Selection Rates
Figure 7: Selection rate as a function of accuracy (left) and FLOPs (right) for different error tolerances on ImageNet1k. The selection rate $\mathbb{P}(r(x)\geq\theta)$ represents the fraction of data handled at a lower cascade tier without deferring to a larger model, based on a threshold $\theta$. Laxer error tolerances (e.g., 5%) yield higher selection rates, as more samples meet the criteria for safe deferral. In contrast, stricter tolerances (e.g., 1%) result in lower selection rates. Both plots illustrate that higher-accuracy or higher-FLOP models generally achieve higher selection rates, especially as the allowable error tolerance increases, reinforcing the stability and practicality of ABC.
@@
The plots in Figure 7 show the selection rate across different accuracy levels and FLOP values on the ImageNet1k dataset. We explore selection rates for three error tolerances, corresponding to maximum allowable error rates of 1%, 3%, and 5%. As shown in the left plot, which compares selection rate versus accuracy of the models used in the ensemble, higher accuracy models tend to achieve higher selection rates, especially when the error tolerance is relaxed (e.g., 5% maximum error). Conversely, stricter error tolerances (e.g., 1%) lead to lower selection rates across all accuracy levels, as fewer samples meet the stringent requirements for safe deferral at lower cascade tiers. In the right plot, which illustrates selection rate versus FLOPs, a similar trend is observed. Higher-FLOP models, which are generally more accurate, are able to safely handle a larger portion of the data at lower tiers, particularly as the error tolerance increases.
@@ D.1 Datasets and Models
Image Tasks ImageNet-1K Deng et al. (2009) Image classification
CIFAR-10 Krizhevsky & Hinton (2009) Image classification
Language Tasks SST-2 Socher et al. (2013) Sentiment analysis
Twitter Financial News Pei et al. (2021) Sentiment analysis
SWAG Zellers et al. (2018) Multiple-Choice QA
Black-Box Experiments GSM8K Cobbe et al. (2021) Math Reasoning
COQA Reddy et al. (2019) Conversational QA
OVERRULING Zheng et al. (2021) Legal Reasoning
HEADLINES Sinha & Khandait (2021) News Classification
@@
Language Models BERT Devlin et al. (2019)
RoBERTa Liu et al. (2019)
XLNet Yang et al. (2019) BASE, LARGE
ELECTRA Radford et al. (2021)
Image Models ResNet He et al. (2016)
ViT Dosovitskiy (2020) Selection based on FLOPs
CLIP Clark et al. (2020)
Black-Box Models LlaMA 3.1 Dubey et al. (2024)
Gemma 2 Team et al. (2024) See Table 1
Qwen 2 Yang et al. (2024)
Table 3: Summary of models used for both benchmark and black-box API experiments, across image and language tasks.
@@ D.2 Details for Black-Box API Experiments
We compare ABC to FrugalGPT Chen et al. (2023), 2 variants of AutoMix Madaan et al. (2023), and MoT LLM Cascade Yue et al. (2024). Although HybridLLM Ding et al. (2024) falls into this category of state-of-the-art cascade methods, it has been shown to underperform FrugalGPT and AutoMix Madaan et al. (2023). For practical evaluation, we implemented all these baseline methods to work in a fully functional cascade system.
@@
  • AutoMix: AutoMix trains a different router for all possible cascade steps, i.e., $n-1$ routers for the n cascade tiers, and this has to be repeated for every new task or model replacement in the system. A threshold (AutoMix+T) or POMDP (AutoMix+P) is trained with a combination of $\geq 50$ training samples from the same test data distribution and the initial inferences generated on the test data by the two models involved in each cascading step. After training the router, using the cascading system often involves running a few-shot self-verification 8 times at a high sampling temperature (temp = 1.0), using the same model that generates inference at the given cascade tier. Automix then averages the self-verification results and meta-verify it with the best parameters of the routing strategy to decide when to exit.
  • FrugalGPT: Just like AutoMix, FrugalGPT needs to train $n-1$ routers for $n$ cascade tiers, and each router needs to have a sense of the data distribution and the model’s predictive power. $\geq 500$ training samples and inference generated on these samples by the tier’s model are needed to train each tier/model’s scorer, a DistilBERT Sanh et al. (2019).
  • MoT LLM Cascade: Yue et al. (2024) focuses on sampling and consistency checking as a means of cascade. To measure consistency, the weaker LLM generates multiple answers for a single question by varying the randomness in the LLM’s responses—i.e., varying the temperature of the model—while using in-context demonstrations and reasoning techniques (e.g., Chain-of-Thought Wei et al. (2022)) to influence how the model generates answers. The system compares the different sampled answers and picks the most consistent one. If the consistency score is high enough, the weaker model’s answer is accepted. Otherwise, the question is passed to the next tier.
@@
We evaluate these methods on a variety of datasets and tasks as shown in Table 2. To ensure consistency in output format and easier evaluation setup, we use few-shot prompting—specifically, 4-shot—for all models across all tasks. For evaluation metrics, we use the macro F1 score for CoQA to capture overlaps between predictions and ground-truth answers, while we measure accuracy (essentially exact match) for the rest of the tasks. In terms of efficiency, we also measure the costs of using the model APIs. It is important to note that AutoMix and FrugalGPT incur additional setup costs that we did not reflect in our reported results. These setup costs and associated latency represent a significant constraint, especially for scenarios requiring frequent retraining or adaptation to new tasks, data distributions, and models.
As shown in Figure 5, we observe that ABC is often more cost-effective than the baselines, even with their more sophisticated routing mechanisms and singular model tiers. Our analyses show that ABC’s efficient deferral strategy allows a more aggressive utilization of cheaper models for a significant portion of the input while only using expensive models when necessary. For instance, we realize—upon analyzing FrugalGPT’s results—that the trained scorer struggles as an efficient deferral mechanism as the tasks get harder; hence, it is more likely to take the safer option to cascade as the difficulty of the test samples increases. This means that ABC can be expected to be more efficient since the scaling laws ensure that the sum of the costs of using several cheaper models is still much less than the cost of using the larger model in the next tier. AutoMix, on the other hand, uses a few-shot self-verification system that is sampled $>1$ times; hence, the additional API calls add significantly to its cost of usage. Considering that the self-verification process is an integral part of the AutoMix setup, it can be guaranteed that ABC will always be cheaper to use than AutoMix, despite using more models.
@@ Appendix E Benefits of ABC
Figure 8: Inference execution in parallel vs non-parallel on CIFAR-10.
Based on Section 5.1.1, we additionally show in Figure 8 the superiority of parallel inference execution for cascading over using the best single model using CIFAR-10 as a case study. We also show that in the worst-case scenarios in which every single inference has to be sequentially produced over each ensemble and cascade, there are still considerable savings over the largest single models.
@@ E.2 Cost Benefits
Table 4: GPU usage costs estimates from Lambda Cloud Lambda (2024), as of September 2024.
Table 5: Table describing various metrics at different cascade tiers for each dataset. Each dataset’s row consists of the following metrics: the fraction of samples processed at each depth, the cost of Total GPU usage per hour at each tier, the average latency of the specific tier (in milliseconds), and the average FLOPs at each tier (floating-point operations). The ABC column for each metric represents the entire cascading system’s average across all cascade depths.

### v3

@@ 6 Conclusion
Although using an ensemble of models can provide a powerful deferral rule for cascading, the additional costs required to compute such an ensemble may not lead to savings in all inference scenarios. Despite this, our work shows that this simple approach is surprisingly effective given the large differences in model sizes that reach state-of-the-art accuracy in recent ML tasks. We demonstrate improvements via a number of real-world case studies, including a study on communication costs in edge-to-cloud inference, rental costs in cloud-based settings, and the cost of black-box API services. Overall, our results demonstrate ABC’s capacity to improve the efficiency of adaptive inference systems without the complexities associated with traditional cascade frameworks, making it a compelling option for practitioners focused on reducing inference latency.
Future Work. Several promising directions emerge from this work. First, extending ABC to open-ended generation tasks would significantly broaden its applicability, particularly given the growing prominence of LLM-based applications. Second, exploring more efficient ensemble methods could further enhance ABC’s benefits in scenarios with limited parallelization capabilities, potentially through techniques that reduce ensemble overhead while maintaining agreement quality. Finally, investigating bias-aware agreement mechanisms or construction strategies that promote diversity in decision-making patterns across demographic groups could address potential bias propagation concerns inherent in majority-based voting systems.
@@ Appendix A Agreement-Based Cascading (ABC) can Improve Accuracy
Proof of Proposition 4.1
Part 1: By Definition 4.1, we have $P(r(x)=0,H_{1}^{k}(x)\neq y)\leq\epsilon$. The risk of the cascade decomposes as:
$R(\mathcal{M}_{r})$ $=P(r(x)=0,H_{1}^{k}(x)\neq y)+P(r(x)=1,h_{2}(x)\neq y)$ (7)
$\leq\epsilon+P(r(x)=1)R(h_{2})\leq R(h_{2})+\epsilon$ (8)
Part 2: The expected cost follows from Equation 1 and conditioning on the deferral decision:
$E[C(\mathcal{M}_{r})]=P(r(x)=0)C(H_{1}^{k})+P(r(x)=1)C(h_{2})=(k^{\rho}\gamma+P(r(x)=1))C(h_{2})$
@@ Appendix B Estimating Agreement Threshold $\theta$
Figure 6: Estimation of agreement threshold $\theta$ stability as a function of the number of samples used, across different model accuracy levels on the ImageNet-1K dataset. Each plot corresponds to a model with a specific accuracy, shown in the legend. The initial estimate is with 100100100 samples, with subsequent estimates with larger and larger number of samples. The initial estimate is reasonably close within subsequent estimates with larger number of samples.
@@ Appendix C Existence of Safe Deferral Rules and Selection Rates
Figure 7: Selection rate as a function of accuracy (left) and FLOPs (right) for different error tolerances on ImageNet-1K. The selection rate $\mathbb{P}(r(x)\geq\theta)$ represents the fraction of data handled at a lower cascade tier without deferring to a larger model, based on a threshold $\theta$. Laxer error tolerances (e.g., 5%) yield higher selection rates, as more samples meet the criteria for safe deferral. In contrast, stricter tolerances (e.g., 1%) result in lower selection rates. Both plots illustrate that higher-accuracy or higher-FLOP models generally achieve higher selection rates, especially as the allowable error tolerance increases, reinforcing the stability and practicality of ABC.
@@
The plots in Figure 7 show the selection rate across different accuracy levels and FLOP values on the ImageNet-1K dataset. We explore selection rates for three error tolerances, corresponding to maximum allowable error rates of 1%, 3%, and 5%. As shown in the left plot, which compares selection rate versus accuracy of the models used in the ensemble, higher accuracy models tend to achieve higher selection rates, especially when the error tolerance is relaxed (e.g., 5% maximum error). Conversely, stricter error tolerances (e.g., 1%) lead to lower selection rates across all accuracy levels, as fewer samples meet the stringent requirements for safe deferral at lower cascade tiers. In the right plot, which illustrates selection rate versus FLOPs, a similar trend is observed. Higher-FLOP models, which are generally more accurate, are able to safely handle a larger portion of the data at lower tiers, particularly as the error tolerance increases.
@@ D.1 Datasets and Models
Image Tasks ImageNet-1K (Deng et al., 2009) Image classification
CIFAR-10 (Krizhevsky & Hinton, 2009) Image classification
Language Tasks SST-2 (Socher et al., 2013) Sentiment analysis
Twitter Financial News (Pei et al., 2021) Sentiment analysis
SWAG (Zellers et al., 2018) Multiple-Choice QA
Black-Box Experiments GSM8K (Cobbe et al., 2021) Math Reasoning
COQA (Reddy et al., 2019) Conversational QA
OVERRULING (Zheng et al., 2021) Legal Reasoning
HEADLINES (Sinha & Khandait, 2021) News Classification
Table 3: Summary of models used for both benchmark and black-box API experiments, across image and text tasks.
@@
Language Models BERT (Devlin et al., 2019)
RoBERTa (Liu et al., 2019)
XLNet (Yang et al., 2019) BASE, LARGE
ELECTRA (Radford et al., 2021)
Image Models ResNet (He et al., 2016)
ViT (Dosovitskiy, 2020) Selection based on FLOPs
CLIP (Clark et al., 2020)
Black-Box Models LlaMA 3.1 (Dubey et al., 2024)
Gemma 2 (Team et al., 2024) See Table 1
Qwen 2 (Yang et al., 2024)
@@ D.2 Details for Black-Box API Experiments
We compare ABC to FrugalGPT (Chen et al., 2023), 2 variants of AutoMix (Madaan et al., 2023), and MoT LLM Cascade (Yue et al., 2024). Although HybridLLM (Ding et al., 2024) falls into this category of SOTA methods, it has been shown to underperform FrugalGPT and AutoMix (Madaan et al., 2023). For practical comparison, we implement all methods in a fully functional cascade system.
@@
  • AutoMix: AutoMix trains a different router for all possible cascade steps, i.e., $n-1$ routers for the n cascade tiers, and this has to be repeated for every new task or model replacement in the system. A threshold (AutoMix+T) or POMDP (AutoMix+P) is trained with a combination of 50$\geq$ 50≥ 50 training samples from the same test data distribution and the initial inferences generated on the test data by the two models involved in each cascading step. After training the router, using the cascading system often involves running a few-shot self-verification 8 times at a high sampling temperature (temp = 1.0), using the same model that generates inference at the given cascade tier. Automix then averages the self-verification results and meta-verify it with the best parameters of the routing strategy to decide when to exit.
  • FrugalGPT: Just like AutoMix, FrugalGPT needs to train $n-1$ routers for $n$ cascade tiers, and each router needs to have a sense of the data distribution and the model’s predictive power. 500$\geq$ 500≥ 500 training samples and inference generated on these samples by the tier’s model are needed to train each tier/model’s scorer, a DistilBERT (Sanh et al., 2019).
  • MoT LLM Cascade: Yue et al. (2024) focuses on sampling and consistency checking as a means of cascade. To measure consistency, the weaker LLM generates multiple answers for a single question by varying the randomness in the LLM’s responses—i.e., varying the temperature of the model—while using in-context demonstrations and reasoning techniques (e.g., Chain-of-Thought (Wei et al., 2022)) to influence how the model generates answers. The system compares the different sampled answers and picks the most consistent one. If the consistency score is high enough, the weaker model’s answer is accepted. Otherwise, the question is passed to the next tier.
@@
We evaluate these methods on a variety of (closed) generation datasets and tasks as shown in Table 2. To ensure consistency in output format and easier evaluation setup, we use few-shot prompting—specifically, 4-shot—for all models across all tasks. For evaluation metrics, we use the macro F1 score for CoQA to capture overlaps between predictions and ground-truth answers, while we measure accuracy (essentially exact match) for the rest of the tasks. In terms of efficiency, we also measure the costs of using the model APIs. It is important to note that AutoMix and FrugalGPT incur extra setup costs that we did not factor into our results. These costs and associated latency represent a significant constraint, especially for scenarios requiring frequent retraining or adaptation to new tasks, distributions, and models.
As shown in Figure 5, we observe that ABC is often more cost-effective than the baselines, even with their sophisticated routing mechanisms and singular model tiers. Our analyses show that ABC’s efficient deferral strategy allows a more aggressive utilization of cheaper models for a significant portion of the input while only using expensive models when necessary. For instance, we realize—upon analyzing FrugalGPT’s results—that the trained scorer struggles as an efficient deferral signal as the tasks get harder; hence, it is more likely to take the safer option to cascade as test sample difficulty increases. This means that ABC can be expected to be more efficient since the scaling laws ensure that the sum of the costs of using several cheaper models is still much less than the cost of using the larger model in the next tier. AutoMix, on the other hand, uses a few-shot self-verification system that is sampled >1>1> 1 times; hence, the additional API calls add significantly to its cost of usage. Considering that the self-verification process is an integral part of the AutoMix setup, it can be guaranteed that ABC will always be cheaper to use than AutoMix, despite using more models.
@@ Appendix E Benefits of ABC
Figure 8: Impact of parallelization on ABC performance for CIFAR-10. Left: With parallel inference execution ($\rho=1$), ABC configurations consistently outperform the best single models across different cascade lengths and ensemble sizes. Right: Even with sequential execution ($\rho=0$), ABC maintains advantages over single models, though with reduced efficiency. The results demonstrate that while parallelization is beneficial, ABC remains effective even under sequential constraints when cost disparities are sufficient.
Based on Section 5.1.1, we additionally show in Figure 8 the superiority of parallel inference execution for cascading over using the best single model using CIFAR-10 as a case study. We also show that in the worst-case scenarios in which every single inference is sequentially produced over each ensemble and cascade, there are still considerable savings over the largest single models, if the scenarios assumptions are met.
@@ E.2 Cost Benefits
Table 4: GPU rental costs from Lambda Cloud (Lambda, 2024) (September 2024) showing the substantial cost disparities between hardware generations. The 25$\times$ cost difference between H100 and V100 GPUs, combined with more modest throughput differences, creates favorable conditions for ABC’s heterogeneous hardware placement strategy described in Section 5.2.2
@@
Table 5: Detailed cost breakdown across cascade tiers for each dataset, showing the fraction of samples processed at each tier, associated GPU costs, latency, and FLOPs. The high fraction of samples processed at cheaper early tiers (52-93%) demonstrates ABC’s effectiveness at concentrating expensive computation on truly difficult samples. ABC consistently outperforms single best models across all metrics while achieving substantial cost savings.